OpenStudy (anonymous):

given the function g(x)=-x^2-4x-3, find g(-1/3)

OpenStudy (luigi0210):

\[\huge \mathbb{\color{orange}{WELCOME ~TO~OPENSTUDY}}\] What you do is plug in -1/3 so solve this: \[\Large g(-\frac{1}{3})=-(\frac{1}{3})^2-4(-\frac{1}{3})-3\]

OpenStudy (anonymous):

How can I solve it though?

OpenStudy (luigi0210):

By evaluating the numbers I provide.

OpenStudy (anonymous):

Is there a way to get rid of the denominators to convert it into normal variables and get the right answer?